# Geometric Origin of Quantum Entanglement

### v1

@@ 5 Conclusion
We have shown that the massless representations of the extended Poincaré group necessarily possess a two–component internal structure correlating forward/backward Wigner sectors. This structure is mathematically equivalent, under local observables, to the entanglement of two qubits. Therefore quantum entanglement has a geometric origin in the additional discrete superluminal symmetry generated by $\Lambda_{\infty}$.
A quantum state tomography experiment doable within standard quantum optics can in principle distinguish the two eigenvalues of $U(\Lambda_{\infty})$ giving experimental evidence of the representation-theoretic structure of massless particles in the extended Poincarè group which are intrinsically entangled objects.
[…]
Figure 1: Schematic of the proposed single–photon experiment. BS1 prepares a superposition of counterpropagating modes $|+\rangle,|$-$\rangle$ (forward/backward sectors). The elements $H,V$ encode polarisation $|H\rangle,|V\rangle$ on each arm, implementing the state $\tfrac{1}{\sqrt{2}}(|+\rangle\otimes|H\rangle+\varepsilon|-\rangle\otimes|V\rangle)$. BS2 measures $\sigma_{x}$ on the direction qubit; in each output, a HWP (Half Wave Plate) + PBS (Polarization Beam Splitter) measures $\sigma_{x}$ on polarisation. Joint detector clicks ($D_{i}$) estimate $\langle\sigma_{x}^{(\text{dir})}\otimes\sigma_{x}^{(\text{pol})}\rangle=\varepsilon$.

### v2

@@ 6 Conclusion
We have studied the massless sector of the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}$ introduced in [1]. Because the lightlike little group is enlarged to $ISO(2)\rtimes\mathbb{Z}_{2}$, the corresponding massless UIRs necessarily carry a binary internal structure: the representation space decomposes as a direct sum of forward and backward Wigner sectors, $\mathcal{H}_{\oplus}=\mathcal{H}_{\mathrm{fwd}}\oplus\mathcal{H}_{\mathrm{bwd}}$, and the discrete involutive element acts as a sector–exchange symmetry. We showed that this binary structure admits an operationally faithful embedding into a bipartite system $\mathcal{H}\otimes\mathbb{C}^{2}$: there exist a sector isometry $V$ and a natural *–homomorphism $\iota$ of observable algebras such that expectation values of all local observables $A\otimes B$ coincide with those of the corresponding operators $\iota(A\otimes B)$ on $\mathcal{H}_{\oplus}$. In a two–mode single–photon restriction, this reduces to ordinary two–qubit Bell-like states and Pauli correlations.
We then connected the abstract label $\varepsilon=\pm 1$ to an experimentally accessible signature. For a fixed propagation axis $\hat{n}$ the two sectors can be realised as counterpropagating optical modes, while a two-level internal degree of freedom (e.g. polarisation) encodes the binary label. In this setting the correlation observable $\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}$ has expectation value $\varepsilon$ on the $U(\Lambda_{\infty})$–eigenstates, providing a direct tomographic handle on the sector degree of freedom and a concrete route to empirical constraints on the extended massless representation.
Finally, we used the same single–photon two–mode restriction as a computational code. The two involution eigenstates define a logical qubit (the single photon entanglement qubit) whose pure states form a Bloch sphere and whose logical generators close $\mathfrak{su}(2)$; physically, a relative phase shifter and a two–mode coupler generate all logical single–qubit unitaries. On two such qubits, a polarisation parity measurement restricts to a logical Pauli–product measurement on the code space. From this primitive we constructed a Choi resource state for $CZ_{L}$ as a stabiliser state and implemented $CZ_{L}$ (hence $CNOT_{L}$) on arbitrary inputs via a gate–teleportation protocol, up to Pauli byproducts. Together, arbitrary single–qubit $SU(2)$ control and an entangling two–qubit gate establish circuit–model universality of the SPE scheme. In the computational model presented, the logical representation space coincides with the physical Hilbert space of the Poincarè group, and logical operations are equivalent to idealized physical transformations. This identification has the consequence that quantum entanglement appears as a structural feature of representation theory rather than an externally engineered resource and the leakage of resources in a computation from computational space to physical degrees of freedom not relevant to the computation is not part of the definition of the computation itself but is in principle restricted to noise and imperfections of optical gates implementing logical operations.
This could be a conceptual breakthrough in the developments of architectures for scalable quantum computers.